White Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates: Department Cutting Finishing Direct labor-hours 6,600 65,000 Machine-hours 57,300 3,200 Total fixed manufacturing overhead cost $ 390,000 $ 443,000 Variable manufacturing overhead per machine-hour $ 3.00 — Variable manufacturing overhead per direct labor-hour — $ 4.75 Required: 1. Compute the predetermined overhead rate for each department. 2. The job cost sheet for Job 203, which was started and completed during the year, showed the following: Department Cutting Finishing Direct labor-hours 3 12 Machine-hours 86 4 Direct materials $ 780 $ 360 Direct labor cost $ 75 $ 300 Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203. 3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates?
Answer #1

1) Predetermine overhead rate :

Cutting department = 390000/57300+3 = 9.81 per machine hour

Finishing department = 443000/65000+4.75 = 11.57 per labor hour

2) Total manufacturing cost

Job 203
Direct material (780+360) 1140
Direct labor 375
Overhead (86*9.81+12*11.57) 982.50
Total manufacturing cost 2497.50 or 2498

3) Yes